OpenStudy (anonymous):

To form 3 adjacent squares, 10 toothpicks are required. How many toothpicks are required to form 12 adjacent squares????

hero (hero):

3/10 =12/x 120 = 3x x = 120/3 x = 40

OpenStudy (anonymous):

n*3+1 12*3+1=37

OpenStudy (anonymous):

i would have said 4 for the first one and then 3 more for each additional one

OpenStudy (anonymous):

so \[4+3\times 11=37\] more than one way to skin a cat

OpenStudy (anonymous):

general formula is 3(n-1) + 4 3(12-1) + 4 = 33 + 4 = 37

hero (hero):

I think I made sure wrong assumptions

OpenStudy (anonymous):